A psychiatrist is testing a new antianxiety drug, which seems to have the potentially harmful side effect of lowering the heart rate. For a sample of 50 medical students whose pulse was measured after 6 weeks of taking the drug, the mean heart rate was 70 beats per minute (bpm). If the mean heart rate for the population is 72 bpm with a standard deviation of 12, can the psychiatrist conclude that the new drug lowers heart rate significantly? (Set the level of significance to 0.01.)

Solution: Step 1: State the hypotheses.

Ho: _______________________________________________________________

Ha: _______________________________________________________________

Step 2: The level of significance and the critical region. 𝛼 = _____, π‘π‘Ÿπ‘–π‘‘π‘–π‘π‘Žπ‘™ π‘£π‘Žπ‘™π‘’π‘’ = _____.

Step 3: Compute for the value of one sample test. π‘π‘œπ‘šπ‘π‘’π‘‘π‘’π‘‘ 𝑑𝑒𝑠𝑑 π‘£π‘Žπ‘™π‘’π‘’ = _______.

Step 4: Decision rule. ____________________________________________________________

Step 5. Conclusion. ______________________________________________________________


1
Expert's answer
2021-10-12T09:39:21-0400

Solution:

"\bar{x}=70 \\

n=50 \\

\mu=72 \\

\sigma = 12"

Step 1: State the hypotheses.

"Ho: \mu = 72 \\

Ha: \mu<72"

Step 2: The level of significance and the critical region. 𝛼

the level of significance = 0.01,

π‘π‘Ÿπ‘–π‘‘π‘–π‘π‘Žπ‘™ π‘£π‘Žπ‘™π‘’π‘’ "Z_{cr} = -2.32."

Step 3: Compute for the value of one sample test.

"Z = \frac{\bar{x} - \mu}{\sigma / \sqrt{n}} \\

Z = \frac{70-72}{12 / \sqrt{50}} = -1.178"

π‘π‘œπ‘šπ‘π‘’π‘‘π‘’π‘‘ 𝑑𝑒𝑠𝑑 π‘£π‘Žπ‘™π‘’π‘’ = -1.178.

Step 4: Decision rule. Reject Ho if "Z \u2264 - Z_{cr}."

Step 5. Conclusion.

"Z=-1.178 > Z_{cr} = -2.32"

Accept Ho. We can conclude, that the new drug does NOT lower heart rate significantly.